OpenStudy (anonymous):

Solve the equation. 2(a − 4) = 3(a + 8)

OpenStudy (anonymous):

Do you know how to start it? :)

OpenStudy (anonymous):

Distribute first.

OpenStudy (anonymous):

Right. SO distribute to ALL #s and then what? :)

OpenStudy (anonymous):

i'm just guessing but i think 2a - 8 = 3a + 24?

OpenStudy (anonymous):

YES now combine like terms...

OpenStudy (anonymous):

2a-8= 3a + 24 2a= 3a + 24 +8 2a= 3a + 32 2a-3a= 32 -a=32 a=-32

OpenStudy (anonymous):

^^ Awesome!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Thanks... :)

OpenStudy (anonymous):

Np :)